6. A sample of 100 fluorescent tubes from the short life Tube Company gives a mean length of life of 20.5 hours with a standard deviation of 1.6 hours. Find a 95% confidence interval for the average length of life of those tubes 


1
Expert's answer
2021-11-25T16:58:42-0500

"n=100 \\

\bar{x} = 20.5 \\

\sigma = 1.6 \\

ME = Z_{α/2} \frac{\sigma}{\sqrt{n}} \\

ME = 1.96 \times \frac{1.6}{\sqrt{100}} = 0.313 \\

Lower \; limit = \bar{x} -ME \\

= 20.5 -0.313 = 20.187 \\

Upper \; limit = \bar{x} + ME \\

= 20.5 + 0.313 = 20.813"